A certain drug is claimed by its manufacturers to reduce overweight men by 4.75 kg per month,

with a standard deviation of 0.89 kg. Ten randomly chosen men reported losing an average of 4.25 kg

within a month. Does this data support the claim of the manufacturer at 0.05 level of significance?


1
Expert's answer
2021-07-05T13:34:31-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=4.75"

"H_1:\mu\not=4.75"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha=0.05," and the critical value for a two-tailed test is "z_c=1.96."

The rejection region for this two-tailed test is "R=\{z:|z|>1.96\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{4.25-4.75}{0.89/\sqrt{10}}\approx -1.77656"

Since it is observed that "|z|=1.77656<1.96=z_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value is "p=2P(Z<-1.77656)=0.075641," and since "p=0.078641<0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population "\mu" is different than "4.75," at the "\alpha=0.05" significance level.